# 4D Scaffold Gaussian Splatting for Efficient Dynamic Scene Reconstruction

### v1

@@ Abstract
Existing 4D Gaussian methods for dynamic scene reconstruction offer high visual fidelity and fast rendering. However, these methods suffer from excessive memory and storage demands, which limits their practical deployment. This paper proposes a 4D anchor-based framework that retains visual quality and rendering speed of 4D Gaussians while significantly reducing storage costs. Our method extends 3D scaffolding to 4D space, and leverages sparse 4D grid-aligned anchors with compressed feature vectors. Each anchor models a set of neural 4D Gaussians, each of which represent a local spatiotemporal region. In addition, we introduce a temporal coverage-aware anchor growing strategy to effectively assign additional anchors to under-reconstructed dynamic regions. Our method adjusts the accumulated gradients based on Gaussians’ temporal coverage, improving reconstruction quality in dynamic regions. To reduce the number of anchors, we further present enhanced formulations of neural 4D Gaussians. These include the neural velocity, and the temporal opacity derived from a generalized Gaussian distribution. Experimental results demonstrate that our method achieves state-of-the-art visual quality and $97.8\%$ storage reduction over 4DGS.
1 Introduction
@@
Reconstructing dynamic scenes from multi-view videos has received significant attention due to its wide applications in augmented reality and virtual reality. Beyond methods based on neural radiance fields (NeRFs) [13, 12, 21, 28, 1, 24], 3D Gaussian Splatting (3DGS) [8] has become a major approach for dynamic scene reconstruction, due to its ability to render high-quality novel views in real-time. The two main categories of this approach are 1) modeling temporal changes as deformations of 3D Gaussians [34, 31, 2] and 2) employing 4D Gaussians to approximate a scene’s 4D volumes [35, 14].
Figure 1: Teaser. Our method achieves superior image quality, high rendering speed, and low storage with single-frame COLMAP [25] points, surpassing other 4D Gaussian methods.
Directly optimizing 4D Gaussians offers faster rendering and higher visual quality than deformation-based methods. Gaussians defined in 4D space can represent sudden changes in the scene, without computing numerous deformations of Gaussians. However, these advantages come at a storage cost: 4D Gaussians demand substantial storage, often exceeding 6GB for 10-second videos. The large number of 4D Gaussians, coupled with extensive parameter counts for 4D properties, hinders scaling this approach for real-world applications. We aim to reduce the high storage demands of 4D Gaussians, by addressing excessively repetitive patterns in dynamic scenes across space and time.
@@
In this paper, we introduce a 4D anchor-based framework for dynamic scene reconstruction, a method that retains the visual quality and the fast rendering speed of 4D Gaussians with significantly reduced storage. In our method, dynamic scenes are represented through structured anchors aligned with a sparse 4D grid. Each anchor holds a compressed feature vector that models a set of neural 4D Gaussians, each representing a local spatiotemporal region. The properties of neural Gaussians are dynamically generated from the anchor feature vector via shared MLPs.
Our method is a non-trivial extension of 3D scaffolding [17] to 4D space, as it should carefully consider how to effectively capture dynamic regions with fewer anchors. Previous 3D scaffolding accumulates gradients of Gaussians to create new anchors, without accounting for temporal coverage. This penalizes dynamic regions that appear shortly, in contrast to background regions that are present in most frames and thus accumulate larger gradients.
@@
We present a temporal coverage-aware anchor growing strategy to allocate new anchors to under-reconstructed dynamic regions. Our approach adjusts the accumulated gradients based on each Gaussian’s temporal coverage: Gaussians with shorter coverage are adjusted to have larger gradients. This modification prevents under-reconstructed dynamic regions from being penalized for their short appearance periods. While the proposed anchor growing operation improves the quality of reconstructed dynamic regions, it also leads to a substantial increase in the number of anchors, and associated storage costs. Such increase originates from the limited capability of generated 4D Gaussians to sufficiently capture real-world dynamics.
To this end, we propose two formulations that enhance our neural Gaussians to better represent actual scene elements. First, we model the time-varying positions of Gaussians through our neural velocity, allowing them to efficiently represent dynamics as a series of piecewise linear movements. Second, we modify our temporal opacity based on a generalized Gaussian function. This modification captures sudden changes in dynamic scenes more effectively, whereas the temporal opacity of previous methods [10, 14] lack this capability. These formulations enhance the capability of Gaussians to capture dynamics in local regions, reducing the number of anchors required for high quality.
Our 4D anchor-based reconstruction pipeline effectively addresses redundancy present in dynamic scenes through compression. As shown in Figure 1, our method achieves state-of-the-art visual quality and rendering speed, while reducing storage demands by $93.8\%$ compared to 4DGS [36].
2 Related work
This section presents a comprehensive review of scene representations, focusing on three aspects. First, we examine approaches for efficient 3D Gaussians (Sec. 2.1). Then, we investigate methods for both deforming 3D Gaussians and utilizing 4D Gaussians in dynamic scene reconstruction (Sec. 2.2). Lastly, we review various approaches employing neural features for scene reconstruction (Sec. 2.3).
@@ 2 Related work
Figure 2: Overview of 4D anchor-based method. We begin with a sparse 4D grid of anchors, each defined by a unique 4D spatiotemporal position $p$ and learnable offsets. Shared MLPs utilize these anchors to generate neural 4D Gaussians, capturing dynamic appearance changes with view direction $\mathbf{d}$. These 4D Gaussians are then projected to 3D Gaussians at specific timesteps for rendering, using a 3D Gaussian splatting pipeline to produce the final rendered image. We omit the $z$-axis for simplicity.
2.1 Efficient 3D Gaussians
@@
In recent years, 3D Gaussian Splatting (3DGS) has attracted significant attention for achieving real-time rendering by representing scenes with 3D Gaussian primitives and introducing a tile-based rasterizer. For photorealistic rendering quality, 3DGS requires a significant number of 3D Gaussians, which increases storage costs. To address this problem, [6, 7, 29] remove unnecessary Gaussians that do not harm rendering quality, [16, 19, 20] replace Gaussians or their parameters with efficient representations, and [19, 4, 33, 18] compress Gaussian parameters using existing compression techniques such as entropy coding and image compression. In this paper, we address eliminating excess Gaussians and improving efficiency in dynamic scenes with 4D Gaussians, which have been relatively unexplored compared to 3D Gaussians.
2.2 Dynamic 3D Gaussians
@@
Two main approaches are proposed to extend 3DGS [8] into dynamic scene reconstruction. The first approach involves deforming 3D Gaussians along with temporal changes [34, 31, 2, 26, 15]. These deformable 3D Gaussians offer the advantage of compact storage requirements but exhibit relatively slow rendering speeds and low visual quality. In contrast, the other approach directly employs 4D Gaussians in the spatio-temporal domain [35, 14]. These approaches demonstrate superior visual quality and faster rendering speeds, but suffer from higher storage requirements. To address this, our method efficiently models temporal opacity and predicts Gaussian properties through neural features.
2.3 Feature-based neural rendering
@@
Recently, a growing trend in scene reconstruction has been to integrate neural features as additional inputs to enhance model performance. For instance, there have been attempts to extract features from source views and utilize them for novel view synthesis, enabling few-view reconstruction [37, 3] or view interpolation through transformers [30, 22, 27]. In 3DGS studies, [5] uses a hash grid instead of per-Gaussian SH coefficients, [14] renders features and decodes them into RGB colors via shallow MLP. [32] generates Gaussian attributes from a multi-level tri-plane. [16] predicts the attributes of local 3D Gaussians from anchor features. Our work focuses on efficient dynamic representations, which are relatively underexplored. We introduce a 4D anchor-based framework that includes dynamic neural velocity and temporal opacity derived from a generalized Gaussian distribution, considering real-world dynamics.
3 Method
@@
We reconstruct dynamic 3D scenes with 4D Gaussians. Especially, our method builds 4D anchors which specify nearby 4D Gaussians (Sec. 3.1). We encourage anchors to be closely placed at under-reconstructed dynamic regions (Sec. 3.2). Furthermore, we design each 4D Gaussian to move along a line segment, and define the temporal opacity function which fits piecewise persistent period (Sec. 3.3).
3.1 Overview
@@
The overview of our method is illustrated in Figure 2. Our method begins with a set of sparse 4D anchor points, each of which has a unique, grid-aligned spatiotemporal 4D position $\mathbf{p}\in\mathbb{R}^{4}$ with a feature vector $\mathbf{f}\in\mathbb{R^{C}}$. We leverage shared MLPs to produce $K$ neural 4D Gaussians from these anchor features. Corresponding 3D Gaussians are computed from these 4D Gaussians to render a frame at timestep $t$. We describe details of our method in the following paragraphs.
@@
Initializing 4D anchors.   Similar with 3D Scaffold-GS [17], we utilize static point clouds to initialize 4D anchor points. We first obtain the static point cloud from multi-view frames at a certain timestep $t$ using Structure-from-Motion (SfM) [25]. The 4D positions of anchors are initialized from a set of voxels $V\subset\mathbb{R}^{N\times 3}$ obtained from this point cloud:
$p_{v}=(x_{v},y_{v},z_{v},t),\ \ \forall v\in V,$ (1)
@@
where ($x_{v}$, $y_{v}$, $z_{v}$) is the center coordinates of the voxel $v$. Each anchor point also accompanies two learnable parameters: a feature vector $\mathbf{f}$, and 4D offsets $\mathbf{o}\in R^{K\times 3}$ for determining positions of $K$ neural 4D Gaussians.
Beginning with 4D anchors initialized with the static point cloud, our method gradually increases anchors to dynamic regions. This is accomplished by our anchor growing strategy, which will be described in Sec. 3.2.
@@
Neural 4D Gaussians.   We leverage shared MLPs and learnable parameters of the 4D anchors to generate $K$ neural 4D Gaussians from each anchor. Specifically, shared MLPs take the anchor feature $\mathbf{f}$ and yield properties of $K$ neural Gaussians, which include time-invariant opacity $\alpha\in\mathbb{R}$, quaternion $\mathbf{q}\in\mathbb{R}^{4}$ and scaling $\mathbf{s}\in\mathbb{R}^{3}$ for the covariance matrix, view-dependent color $\mathbf{c}\in\mathbb{R}^{3}$. Our shared MLPs also produce temporal scale $\sigma\in\mathbb{R}$ to compute our temporal opacity, and the neural velocity $\mathbf{v}\in\mathbb{R}^{3}$. Note that the color MLP additionally takes the view direction $\mathbf{d}\in\mathbb{R}^{3}$ as inputs for view-dependent modeling.
@@
Rendering neural Gaussians.   To render the neural 4D Gaussians, we compute 3D Gaussian parameters at time $t$ from our time-varying properties. For the $k$-th Gaussian of the anchor $\mathbf{p}$, the opacity $\alpha_{k}(t)$ and the center $\mu_{k}(t)$ of the corresponding 3D Gaussian at time $t$ are derived as
$\mu_{k}(t)=\mu_{k,1:3}+h(t,\mu_{k,4},\mathbf{v}),$ (2)
$\alpha_{k}(t)=\alpha_{k}\cdot g(t,\mu_{k,4},\sigma_{k}),$ (3)
@@
where $\mu_{k}=\mathbf{p}+\mathbf{o}_{k}$ is the coordinates of the 4D Gaussian. $h(\cdot)$, $g(\cdot)$ model time-varying values of the positions and the opacity with the neural velocity and temporal opacity function, which will be further described in Sec. 3.3.
@@
After deriving 3D Gaussians at time $t$, we utilize the existing efficient 3D Gaussian splatting pipeline [8] to render the frames. Same as 3D Scaffold-GS [17], Gaussians within the view frustum and having opacity higher than the threshold are passed to the renderer.
Figure 3: Time-variant Gaussian position calculation. The 3D Gaussian position at rendering time is computed as the sum of the anchor position, the 4D Gaussian offset, and the temporal linear movement. The temporal linear movement is the product of the neural velocity and the time difference, allowing the Gaussian to traverse 3D space over time.
3.2 Temporal coverage-aware anchor growing
@@
To this end, we propose a temporal coverage-aware anchor growing operation. The core of our anchor growing is the computation of the accumulated gradients of each 4D Gaussian $\nabla_{g}$. We formulate $\nabla_{g}$ as a weighted sum of the gradient over $N$ iterations:
$\nabla_{g}=\frac{\sum^{N}w(t,\sigma)\|\nabla_{\text{2D}}\|}{\sum^{N}w(t,\sigma)},$ (4)
where $\nabla_{\text{2D}}$ is the 2D position gradient of the Gaussian. We define the weight term $w$ as a function of $\alpha(t)$ and $\sigma$, with $\gamma$ as a hyperparameter:
$w(t,\sigma)=\alpha(t)(1/\sigma)^{\gamma}.$ (5)
We collect the accumulated gradients of all Gaussians and then voxelize them. New anchors are placed at the centers of the voxels having gradients higher than a threshold.
In contrast, the previous anchor growing strategy [17] designed for static scenes gathers the gradients as a mean over $N$ iterations:
@@
The first difference of our approach compared to the previous anchor growing strategy [17] is that the gradients are accumulated only when the Gaussian is activated. Our approach accurately gathers the gradients of 4D Gaussians placed at the regions where dynamic scene elements only appear in a short period. The second difference is that the temporal coverage $\sigma$ directly influences to the anchor growing operation. This makes regions having a short temporal coverage easier to grow. By modifying the accumulated gradients as a weighted sum, our approach effectively places new anchors to under-reconstructed dynamic regions.
3.3 Modeling 4D Gaussians for compact anchors
To represent dynamic scenes with fewer anchors, each neural Gaussian should be able to model the scene element in a sufficient time range. This necessitates neural 4D Gaussians to accurately model the real-world dynamics.
We introduce two key formulations of 4D Gaussians that improve the capability of Gaussians of modeling real-world dynamics. We model local movements through the neural velocity, and we modify the temporal opacity based on a generalized Gaussian distribution.
Neural velocity.   In computer graphics, one common way to model the motion of an object is through a set of piecewise linear segments. Based on this concept, we propose the neural velocity to model the motions of scene elements with a set of 4D Gaussians having linear movements (see Figure 3). As described in Equation (2), the time-varying 3D position of the 3D Gaussian $\mu_{k}(t)\in\mathbb{R}^{3}$ is determined by the time-varying function $h(\cdot)$. We formulate the $h(\cdot)$ as the linear movement based on the neural velocity $\mathbf{v}$:
$h(t,\mu_{k,4},\mathbf{v})=(t-\mu_{k,4})\mathbf{v}.$ (7)
Note that our neural velocity simplifies the temporal slicing of 4D positions in 4DGS [10], without requiring excessive parameters for 4D covariances.
Despite its simplicity, our neural velocity enables each 4D Gaussian to cover dynamic movements of the elements in a sufficient time range. Such capability reduces the number of neural 4D Gaussians and anchors as a result.
Figure 4: Illustration of the modified temporal opacity. To reconstruct the sudden and continuous appearance of an object, the previous univariate Gaussian model requires the sum of multiple Gaussians, while the generalized Gaussian achieves the same expressiveness with only a single Gaussian.
Modified temporal opacity.   Previous 4D Gaussian methods [10, 14] adopt the univariate Gaussian function as temporal opacity to make Gaussians appear within a specific time range. However, this univariate Gaussian lacks the ability to model sudden changes in real-world scene elements, resulting in excessive Gaussians to model a single scene element. This issue is also discussed in the concurrent work [11], which proposes a simplified Gaussian mixture model as a solution.
Therefore, we propose to modify the temporal opacity function to remove such redundant Gaussians (see Figure 4). We reformulate the temporal opacity function based on a generalized Gaussian distribution. This offers steeper derivatives at the beginning and the end, making it better suited to fit piecewise persistent periods. Our time-varying opacity $g(\cdot)$ in Equation (3) is formulated as
$g(t,\mu_{k,4},\sigma_{k})=e^{-(|t-\mu_{k,4}|/\sigma_{k})^{\beta}}.$ (8)
This formulation greatly reduces the number of 4D Gaussians and anchors to represent dynamic scenes while preserving the visual quality. In practice, we set the inverse sigma as a learnable parameter, as it leads to more stable training. We also model $\beta=2\beta^{\prime}$ and $\beta^{\prime}$ as a hyperparameter, which removes the ||$\cdot$|| ⋅ | in the above equation.
3.4 Training and implementation details
Table 1: Quantitative results on N3DV dataset. 1 only includes the flame_salmon for storage and FPS [2]. 2 divides the scene into 50 frames and trains each using a separate model. 3 uses COLMAP [25] for all 300 frames.

### v2

# 4D Scaffold Gaussian Splatting with Dynamic-Aware Anchor Growing for Efficient and High-Fidelity Dynamic Scene Reconstruction

@@ Abstract
Modeling dynamic scenes through 4D Gaussians offers high visual fidelity and fast rendering speeds, but comes with significant storage overhead. Recent approaches mitigate this cost by aggressively reducing the number of Gaussians. However, this inevitably removes Gaussians essential for high-quality rendering, leading to severe degradation in dynamic regions. In this paper, we introduce a novel 4D anchor-based framework that tackles the storage cost in different perspective. Rather than reducing the number of Gaussians, our method retains a sufficient quantity to accurately model dynamic contents, while compressing them into compact, grid-aligned 4D anchor features. Each anchor is processed by an MLP to spawn a set of neural 4D Gaussians, which represent a local spatiotemporal region. We design these neural 4D Gaussians to capture temporal changes with minimal parameters, making them well-suited for the MLP-based spawning. Moreover, we introduce a dynamic-aware anchor growing strategy to effectively assign additional anchors to under-reconstructed dynamic regions. Our method adjusts the accumulated gradients with Gaussians’ temporal coverage, significantly improving reconstruction quality in dynamic regions. Experimental results highlight that our method achieves state-of-the-art visual quality in dynamic regions, outperforming all baselines by a large margin with practical storage costs.
Figure 1: Dynamic region quality comparison. (left) Comparison of dynamic region quality ($y$-axis) versus storage cost ($x$-axis) on N3DV. (right) Results on the flame_salmon scene from N3DV. Scaff-naive refers to the naive anchor-based model without our neural Gaussian design and dynamic-aware anchor growing. Our method significantly outperforms all baselines with efficient storage usage, while other methods either exhibit degraded quality in dynamic regions or excessive storage costs.
Introduction
@@
Reconstructing dynamic scenes from multi-view videos has received significant attention due to its wide applications. Beyond methods based on neural radiance fields (NeRFs) (Li et al. 2022; Wang et al. 2022; Attal et al. 2023), 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) has become a major approach for dynamic scene reconstruction, due to its ability to render high-quality novel views in real-time. The two main categories of this approach are 1) modeling temporal changes as deformations of canonical 3D Gaussians (Yang et al. 2023; Bae et al. 2024; Jiawei et al. 2024) and 2) employing 4D Gaussians to approximate a scene’s 4D volumes (Yang et al. 2024; Li et al. 2023; Lee et al. 2024c, a).
Directly optimizing 4D Gaussians offers higher visual quality and faster rendering than deformation-based methods. By modeling temporal changes through multiple 4D Gaussians, each covering a certain time range, they effectively capture complex dynamic regions without the heavy computation cost of deformation fields. However, this approach accompanies a large number of 4D Gaussians, which subsequently leads to substantial storage overhead–often exceeding 6GB for a 10-second video (see Figure 1, left).
Several approaches have attempted to address the storage cost by reducing the number of Gaussians. They achieve this by more complicated motion modeling (Lee et al. 2024b), aggressive pruning (Li et al. 2023), or motion interpolation (Lee et al. 2024a). While these methods effectively reduce storage, they also inevitably remove Gaussians in dynamic regions, sacrificing the expressiveness needed to represent complex temporal changes. As a result, these methods often suffer from quality degradation in dynamic components.
@@
In this paper, we introduce a 4D anchor-based framework that addresses the storage overhead from a different perspective. Instead of reducing the number of Gaussians–which is crucial for maintaining the expressiveness–our method maintains a sufficient number to render dynamic regions with high visual quality. This is achieved by representing dynamic scenes through structured anchor features (Lu et al. 2024a), aligned with a sparse 4D grid. Each anchor holds a compact feature vector to model a local spatiotemporal region. This feature is processed by shared multi-layer perceptrons (MLP) to output properties of neural Gaussians.
Considering the limited capacity of the shared shallow MLP, we propose a compact parametrization of the neural Gaussians that effectively captures temporal changes. This includes two key modelings: linear motion to model the time-varying 3D position, and a generalized Gaussian function to model temporal opacity, which is better suited for capturing sudden appearance changes. This design enables our framework to effectively represent complex dynamic trajectories through a sequence of linear segments.
While the anchor-based scheme effectively reduces storage, a naïve extension of 3D scaffolding (Lu et al. 2024a) fails to accurately capture dynamic regions (see Figure 1, Scaff-naive). The limitation arises from the anchor growing strategy used in (Lu et al. 2024a), which is originally designed for static scenes. In our 4D framework, each anchor and its neural Gaussians are active only within a specific time range, while the anchor growing strategy accumulates gradients across all frames without considering this coverage. As a result, dynamic regions appearing in only a few frames receive less gradient signals, leading to under-reconstructed dynamic components in the final scene.
@@
We present a dynamic-aware anchor growing strategy to properly allocate new anchors to under-reconstructed dynamic regions. Our approach adjusts the accumulated gradients based on each Gaussian’s temporal coverage. We adjust dynamic Gaussians with shorter coverage to have larger gradient, compensating for the penalty caused by their short appearance. This modification encourages new anchors to be allocated more in under-reconstructed dynamic regions, achieving substantial improvement in reconstruction quality.
Our anchor-based framework, coupled with the dynamic-aware anchor growing strategy, represents complex dynamic components in high-quality while addressing storage costs. We validate our approach through extensive experiments conducted on the N3DV (Li et al. 2022) and Technicolor (Sabater et al. 2017) datasets. Notably, our method outperforms all baselines by a large margin with practical storage overhead, as illustrated in Figure 1.
Related Work
@@
Figure 2: Proposed 4D anchor-based framework. We begin with a sparse 4D grid of anchors, each defined by a unique 4D spatiotemporal position $\mathbf{p}$ and learnable offsets. Shared MLPs utilize these anchors to generate neural 4D Gaussians, capturing dynamic appearance changes with view direction $\mathbf{d}$. These 4D Gaussians are then projected to 3D Gaussians at specific time $t_{r}$ for rendering, using a 3D Gaussian splatting pipeline to produce the final rendered image. We omit the $z$-axis for simplicity.
@@
Efficient 3D Gaussians.  In recent years, 3D Gaussian Splatting (3DGS) has attracted significant attention for achieving real-time rendering by representing scenes with 3D Gaussian primitives and introducing a tile-based rasterizer. For photorealistic rendering quality, 3DGS requires a significant number of 3D Gaussians, which increases storage costs. To address this problem, previous works remove unnecessary Gaussians that do not harm rendering quality (Fan et al. 2023; Girish, Gupta, and Shrivastava 2023), replace Gaussians or their parameters with efficient representations (Niedermayr, Stumpfegger, and Westermann 2024; Papantonakis et al. 2024), and compress Gaussian parameters using existing compression techniques such as entropy coding and image compression (Niedermayr, Stumpfegger, and Westermann 2024; Chen et al. 2025; Xie et al. 2025). In this paper, we address compressing Gaussians in spatiotemporal space, which have been relatively unexplored compared to 3D Gaussians.
@@
Dynamic 3D Gaussians.  Two main approaches are proposed to extend 3DGS (Kerbl et al. 2023) into dynamic scene reconstruction. The first approach involves deforming 3D Gaussians along with temporal changes (Yang et al. 2023; Wu et al. 2023; Bae et al. 2024; Shaw et al. 2024; Jiawei et al. 2024; Kwak et al. 2025). These deformable 3D Gaussians offer the advantage of compact storage requirements but exhibit relatively slow rendering speeds and low visual quality. In contrast, the other approach directly employs 4D Gaussians in the spatio-temporal domain. 4DGS (Yang et al. 2024) demonstrate superior visual quality and faster rendering speeds, but suffer from higher storage requirements. Although some works (Li et al. 2023; Lee et al. 2024a, c) improve storage efficiency using fewer Gaussians, they tend to focus on the holistic scene, thereby neglecting the quality of dynamic areas. To address this, we take an alternative approach: reducing storage while maintaining quality by preserving the Gaussian count.
@@
Feature-based neural rendering.  Recently, a growing trend in scene reconstruction has been to integrate neural features as additional inputs to enhance model performance. For instance, there have been attempts to extract features from source views and utilize them for novel view synthesis, enabling few-view reconstruction (Yu et al. 2021; Chen et al. 2021) or view interpolation through transformers (Wang et al. 2021; Reizenstein et al. 2021; T et al. 2023). In 3DGS studies, Compact3DGS (Deng et al. 2024) uses a hash grid instead of per-Gaussian SH coefficients and STG (Li et al. 2023) renders features into RGB via shallow MLPs. Some works generate Gaussian attributes from a multi-level tri-plane (Wu and Tuytelaars 2024) and predict the attributes of local 3D Gaussians from anchor features (Lu et al. 2024a). We introduce a 4D anchor-based framework that includes dynamic linear motion and temporal opacity derived from a generalized Gaussian distribution, considering real-world dynamics.
Method
@@
We reconstruct dynamic 3D scenes through a sparse, grid-aligned 4D anchor grid. Each anchor holds a compressed feature vector that specifies nearby 4D Gaussians. The dynamic-aware anchor growing strategy encourages new anchors to be closely placed at under-reconstructed dynamic regions. Furthermore, we design each 4D Gaussian to move along a line segment, and define the temporal opacity function which fits a piecewise persistent period. Through these components, our framework employs a sufficient number of Gaussians to achieve high-quality reconstruction of dynamic regions, while the storage overhead is significantly reduced via compressed anchor features.
4D Anchor-Based Framework
@@
The overview of our method is illustrated in Figure 2. Our method begins with a set of sparse 4D anchor points, each of which has a unique, grid-aligned spatiotemporal 4D position $\mathbf{p}\in\mathbb{R}^{4}$ with a feature vector $\mathbf{f}\in\mathbb{R}^{C}$. We leverage shared MLPs to produce $K$ neural 4D Gaussians from these anchor features. Corresponding 3D Gaussians are computed from these 4D Gaussians to render a frame at timestep $t$.
@@
Initializing 4D anchors.  Similar to 3D Scaffold-GS (Lu et al. 2024b), we utilize static point clouds to initialize 4D anchor points. We first obtain the static point cloud from multi-view frames at a certain timestep $t_{0}$ using Structure-from-Motion (SfM) (Schönberger and Frahm 2016). The 4D positions of anchors are initialized from a set of voxels $\mathbf{V}\in\mathbb{R}^{N\times 3}$ obtained from this point cloud:
$\mathbf{p}=(x_{v},y_{v},z_{v},t_{0}),\ \ \forall v\in\mathbf{V},$ (1)
$\mathbf{x}_{k}=\mathbf{p}+\Delta\mathbf{x}_{k},\ \ k\in\{1,2,...,K\},$ (2)
@@
where ($x_{v}$, $y_{v}$, $z_{v}$) is the center coordinates of the voxel $v$. Each anchor point also accompanies two learnable parameters: a feature vector $\mathbf{f}$, and 4D offsets $\Delta\mathbf{x}_{k}\in\mathbb{R}^{4}$ for determining properties of $K$ neural 4D Gaussians. 4D Gaussian position $\mathbf{x}$ is determined by anchor position and offset.
@@
Neural 4D Gaussians.  We leverage shared MLPs and learnable parameters of the 4D anchors to generate $K$ neural 4D Gaussians from each anchor. Specifically, shared MLPs take the anchor feature $\mathbf{f}$ and yield properties of $K$ neural Gaussians, which include base opacity $\rho\in\mathbb{R}$, quaternion $\mathbf{q}\in\mathbb{R}^{4}$ and scaling $\mathbf{s}\in\mathbb{R}^{3}$ for the covariance matrix, view-dependent color $\mathbf{c}\in\mathbb{R}^{3}$. Our shared MLPs also produce a temporal scale $\sigma\in\mathbb{R}$ to compute our temporal opacity, and the neural velocity $\mathbf{u}\in\mathbb{R}^{3}$. Note that the color MLP additionally takes the view direction $\mathbf{d}\in\mathbb{R}^{3}$ as inputs for view-dependent modeling.
@@
Rendering neural Gaussians.  To render the neural 4D Gaussians, we compute 3D Gaussian parameters at time $t_{r}$ from our time-varying properties. For the $k$-th Gaussian of the anchor $\mathbf{p}$, the center $\mu_{k}\in\mathbb{R}^{3}$ and the opacity $\alpha_{k}\in\mathbb{R}$ of the corresponding 3D Gaussian at time $t_{r}$ are derived as
$\mu_{k}=\mathbf{x}^{xyz}_{k}+h(t_{r},\mathbf{x}^{t}_{k},\mathbf{u}),$ (3)
$\alpha_{k}=\rho_{k}\cdot g(t_{r},\mathbf{x}^{t}_{k},\sigma_{k}),$ (4)
@@
where $\mathbf{x}^{xyz}_{k}\in\mathbb{R}^{3}$ and $\mathbf{x}^{t}_{k}\in\mathbb{R}$ are the spatial and the temporal position of the 4D Gaussian. $h(\cdot)$, $g(\cdot)$ model time-varying values of the positions and the opacity with the neural velocity and temporal opacity function, which will be further described in the following section.
@@
After deriving 3D Gaussians at time $t_{r}$, we utilize the existing 3D Gaussian splatting pipeline (Kerbl et al. 2023) to render the frames. Same as 3D Scaffold-GS (Lu et al. 2024b), only Gaussians within the view frustum and having opacity higher than the threshold are passed to the renderer.
Compact Parametrization of 4D Gaussians
The properties of neural 4D Gaussians are predicted by a shallow shared MLP. Increasing the number of Gaussian properties complicates the optimization process. To this end, we propose a compact parametrization of 4D Gaussians that captures temporal changes as a set of linear segments. Our neural Gaussian design involves two key modelings. First, we represent time-varying spatial positions as linear motions. Second, we employ a generalized Gaussian function to model temporal opacity, which effectively captures sudden appearance changes with a single parameter.
Linear motion.  As described in Eq. (3), the time-varying 3D position of the 3D Gaussian $\mu_{k}\in\mathbb{R}^{3}$ is determined by the time-varying function $h(\cdot)$. We formulate the $h(\cdot)$ as a linear motion based on the neural velocity $\mathbf{u}$:
$h(t,\mathbf{x}^{t}_{k},\mathbf{u})=(t-\mathbf{x}^{t}_{k})\mathbf{u}.$ (5)
This formulation simplifies the temporal slicing of Gaussian positions proposed in 4DGS (Lee et al. 2024a), using only 3 parameters per Gaussian. Despite its simplicity, a set of Gaussians with linear motion effectively captures the temporal dynamics of scene elements, particularly when combined with our modified temporal opacity.
Figure 3: Illustration of the modified temporal opacity. To reconstruct the sudden and continuous appearance of an object, the univariate Gaussian function requires the sum of multiple Gaussians, while the generalized Gaussian achieves the same expressiveness with only a single Gaussian.
Modified temporal opacity.  Previous 4D Gaussian methods (Yang et al. 2024; Li et al. 2023) typically adopt the univariate Gaussian function to make Gaussians appear within a specific time range. However, this formulation lacks the capacity to model abrupt changes in real-world scene elements, limiting the expressiveness of individual Gaussians.
We reformulate the temporal opacity function based on a generalized Gaussian distribution. This offers steeper derivatives at the beginning and the end, making it better suited to fit piecewise persistent periods. Our time-varying opacity $g(\cdot)$ described in Eq. (4) is formulated as
$g(t,\mathbf{x}^{t}_{k},\sigma_{k})=exp({-(|t-\mathbf{x}^{t}_{k}|/\sigma_{k})^{\beta}}),$ (6)
where $\beta$ is a tunable hyperparameter. In practice, we set the inverse sigma as a learnable parameter, as it leads to more stable training. We also model $\beta=2\beta^{\prime}$ and $\beta^{\prime}$ as a hyperparameter, which removes the $|\cdot|$ in the above equation.
While the issue of temporal opacity is also explored in (Lee et al. 2024a), we show that efficiently representing temporal coverage with fewer parameters is crucial for the performance of the anchor-based scheme.
Dynamic-Aware Anchor Growing
@@
The previous anchor growing strategy (Lu et al. 2024b) designed for static scenes gathers the gradients as a mean over $N$ iterations:
@@
To address this, we propose a dynamic-aware anchor growing operation. The core of our anchor growing is the computation of the accumulated gradients of each 4D Gaussian $\nabla_{g}$. We formulate $\nabla_{g}$ as a weighted sum of the gradient over $N$ iterations:
$\nabla_{g}=\frac{\sum^{N}w(\alpha^{\prime},\sigma)\|\nabla_{\text{2D}}\|}{\sum^{N}w(\alpha^{\prime},\sigma)},$ (8)
where $\nabla_{\text{2D}}$ is the 2D position gradient of the Gaussian, and $\alpha^{\prime}=g(t_{r},\mathbf{x}^{t},\sigma)$ is the time-variant component of the Gaussian opacity. We define the weight term $w$ as a function of $\alpha$ and $\sigma$, with $\gamma$ as a hyperparameter:
$w(\alpha^{\prime},\sigma)=\alpha^{\prime}(1/\sigma)^{\gamma}.$ (9)
We collect the accumulated gradients of all Gaussians and then voxelize them. New anchors are placed at the centers of the voxels having $\nabla_{\text{2D}}$ higher than each threshold.
The dynamic-aware anchor growing strategy differs from the original anchor growing operation (Lu et al. 2024a) in two key aspects. First, gradients are accumulated only when the Gaussian’s temporal opacity is greater than zero (i.e., when it is activated). It allows our method to accurately gather the gradients of 4D Gaussians placed at the regions where dynamic scene elements only appear in a short period. Second, the temporal coverage $\sigma$ directly influences to the anchor growing operation. This encourages regions with a short temporal coverage to receive stronger gradients, making it easier for anchors to grow in those areas. By accumulating gradients as a weighted sum, our method effectively places new anchors to under-reconstructed dynamic regions.
Training and Implementation Details
Dynamic region Full region Computational cost